OpenStudy (anonymous):

The length of an isosceles triangle is increasing at a rate of 5 m/s. At what rate is the hypotenuse changing?

OpenStudy (anonymous):

h= sqrt(2a^2)

OpenStudy (anonymous):

I can implicitly drive the above equation but I don't think that is what they want.

OpenStudy (anonymous):

(1/2)(2a^2)^(-1/2) * 4a = dh/da

OpenStudy (anonymous):

What equation relates the length of the triangle to its hypotenuse?

OpenStudy (anonymous):

like...?

OpenStudy (anonymous):

(sqrt(2a^2))^2/2 = area of triangle

OpenStudy (ranga):



OpenStudy (anonymous):

I was thinking that it was a right triangle.

OpenStudy (ranga):

I think the word "hypotenuse" implies a right triangle which also happens to be isosceles in this problem.

OpenStudy (anonymous):

ya it and isosceles

OpenStudy (anonymous):

an*

OpenStudy (ranga):

@znimon: Can you look at the diagram above and figure out what h is?

OpenStudy (anonymous):

sqrt2

OpenStudy (ranga):

yes \[\sqrt{2}\] So if the side is of length s, then \[h = \sqrt{2}s\] If s changes at the rate of 5m/s, then what is the rate of change of h?

OpenStudy (anonymous):

sqrt(2)*5

OpenStudy (ranga):

Yes. \[5\sqrt{2}\]

OpenStudy (anonymous):

So you get to assume it is a right triangel?

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

I figured I'd make a graphic that's oriented differently than the previous one.

OpenStudy (anonymous):

I got that the rate of change with respect to time for the hypotenuse is 5.

OpenStudy (ranga):

@wio: Is the word hypotenuse used in a context other than a right triangle?

OpenStudy (wolf1728):

The again what is the length of an isosceles triangle - the height? the length of one side?

OpenStudy (ranga):

The length in this context refers to the two equal sides of the triangle with the third side being the hypotenuse. This is a special case of a right triangle with two equal sides and hence the words hypotenuse and isosceles.

OpenStudy (wolf1728):

Then it is definitely a 45 45 90 triangle right?

OpenStudy (ranga):

Yes.

OpenStudy (wolf1728):

It seems to me (according to my graphic) that as the length (or height) of the triangle increases, the hypotenuse increases by twice that amount.

OpenStudy (ranga):

If you assume the words "length of an isosceles triangle" refers to the height of the triangle then you are right. But my interpretation is the length refers to the two equal sides of the isosceles triangle and the hypotenuse refers to the third side of the triangle. In the word problems I have seen, I have come across the phrase "the lengths of the sides of a triangle are a, b & c". But I have not come across the height of a triangle being referred to as the length of a triangle so far.